OpenStudy (anonymous):

differentiate f(x)= 1/2x^6+3x^4-3/4x^3+1/3x^2-7/x+x-4

OpenStudy (anonymous):

hehe now this one looks fun.

OpenStudy (anonymous):

its hell

OpenStudy (anonymous):

break it up into pieces using the sum rule

OpenStudy (anonymous):

so lets do this, shall we ?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

so looking at : http://www.allaboutcircuits.com/vol_5/chpt_6/6.html we want to break it up using the sum rule, so how would that look to you ?

OpenStudy (anonymous):

and then once you have that broken up.. you break it up even further :)

OpenStudy (anonymous):

o0

OpenStudy (anonymous):

you got it ?

OpenStudy (anonymous):

3x^5+12x^3-21/4x^2+

OpenStudy (anonymous):

then the x in the denominator is like 0o

OpenStudy (anonymous):

is it fine so far?

OpenStudy (anonymous):

hmm let me see

OpenStudy (anonymous):

taking the derivitive of the whole equation

OpenStudy (anonymous):

derivative*

OpenStudy (anonymous):

1/2x^6+3x^4-3/4x^3+1/3x^2-7/x+x-4

OpenStudy (anonymous):

so we start wtih this now if we break it up ..

OpenStudy (anonymous):

dx(1/2^6) + dx(3x^4) - dx(3/4x^3) + dx(1/3x^2) - dx(7/x) +dx(x) - dx(4)

OpenStudy (anonymous):

what is the derivative of a constant? 0 because it never changes so we can eliminate that

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so simplify the simple ones first

OpenStudy (anonymous):

1/2x^6 = 3x^5

OpenStudy (anonymous):

we know the dx(x) = x

OpenStudy (anonymous):

so now we have: dx(1/2^6) + dx(3x^4) - dx(3/4x^3) + dx(1/3x^2) - dx(7/x) + x

OpenStudy (anonymous):

3x^5+12x^3-2(1/4)x^2

OpenStudy (anonymous):

left side follows this order: power rule, power rule, power rule, power rule, now sure of dx(7/x)

OpenStudy (anonymous):

then the denominator with 3x^2

OpenStudy (anonymous):

just leave it was 1/6x?

OpenStudy (anonymous):

notsure of dx(7/x)

OpenStudy (anonymous):

would that just be 7?

OpenStudy (anonymous):

-7x^2

OpenStudy (anonymous):

I hope this helps?

OpenStudy (anonymous):

yea, put me in the right direction, thanks

OpenStudy (anonymous):

oh wait dx (7/x ) uses the quotient rule using the quotient rule: f(x) = 7 g(x) = x x * dx(7) - 7 * dx(x) / (x^2)

OpenStudy (anonymous):

which simplfies into -7x / x^2

OpenStudy (anonymous):

ooh

OpenStudy (anonymous):

thanks man big help

OpenStudy (anonymous):

great to hear.

OpenStudy (anonymous):

\[\frac{x^6}{2}+3 x^4-\frac{3 x^3}{4}+\frac{x^2}{3}+x-\frac{7}{x}-4 \]The derivative of the above is:\[3 x^5+12 x^3-\frac{9 x^2}{4}+\frac{7}{x^2}+\frac{2 x}{3}+1 \]